Engineering/D-Night 20180516 (UT) - 2nd half Handover 00:16 AZ time

Night Plan:

  • D-time observations at start
    • DONE Olga - Q1330+011 Binocular MODS (MODS1-Dual MODS-2 Dual) ~ 1 hour
  • AO/ESM/SL comparison of the same target ~ 2.0-2.5 hours
    • We have data from a week ago with natural seeing of ~1.7\x94 (preliminary results look good)
    • Collect ESM data with natural seeing of 0.8" < Natural Seeing < 1.2"
    • Should collect three sets (15 min each)
      • DONE ESM with a \x93bright\x94 AO Reference star: R=9.4
      • DONE Seeing limited data, for comparison
      • DONE ESM with a \x93dim\x94 AO Reference star: R=15.1
      • DONE Full AO performance with the N3.75 to get correction as a function of radial distance from the AO Reference star (AO Ref R=9.4) - caveat is that it was close to sunset - data quality check underway
    • X Near-IR observations of Mergers - end of night (Barry) - I gave up the time to help AO.....


OSA: G. Bechetti

ISp: A. Cavallaro

ISA: B. Rothberg

AO: D. Miller

Mtn Manager: P. Hartley

Observers: B. Rothberg, O. Kuhn

=====================================

UT 07:23 - handed over from AZ - reconfiguring for MODS - this was much later than supposed to be - AZ observers sent telescope across sky from current field (likely loaded the wrong co-pointing/collimation script from target earlier in the evening), then loaded another field and continued to take data past 00:16.

UT 07:50 - Pointing

UT 07:51 - Collimating

UT 07:52 - Moving to QSO (O. Kuhn)

UT 07:58 - Acquiring on both MODS, offset +10" DETY on MODS2 to avoid bad quadrant, AGWs show ~0.9-1" seeing

UT 08:04 - Starting spectroscopic observations of QSO MODS-1 0005-B/0009-R ; MODS-2 0007-B/0017-R - MODS is (almost) FULLY operational! (paraphrasing Olga)

UT 08:25 - 0.8" WFS AGWs - DIMM 0.66"

UT 08:41 - Completed 3 exposures per channel per MODS - taking thru-slit image - MODS-1 looked a few pixels offset to the right, MODS-2 spot-on

UT 08:45 - moving to spectro-photometric standard Hz44

UT 08:51 - MODS2-Red shows electronic noise, went from Elev 32 to Elev 44 Rotator ~ 103 MODS-2

UT 08:58 - taking specphot std data

================================================================

UT 09:14 - Switching to LUCI-BINO for ESM/AO tests with M92

UT 09:42 - Sending preset to M92 for ESM

UT 09:45 - LUCI-1 filter error FW2 - had to initialize in the ENG GUI via LMS

UT 09:49 - still trying to close loops - preset failed

UT 09:51 - resent preset - attempting to close loops

UT 09:56 - SX loop closed, awaiting DX loop

UT 09:59 - DX loops closed - proceeding with script

UT 10:02 - first image shows ~0.27"-0.3" FWHM on LUCI-1

M92 LUCI-1#0003-0013 LUCI-2 #0368-0378 (ESM R=9.43)

UT 10:18 - ESM script done, sending SL script

M92 LUCI-1 #0014-24 LUCI-2 #0379-0389

UT 10:37 - SL Script done, sending dim ESM script (R=15.1)

UT 10:49 - ESM poor continuing

M92 LUCI-1 #0025-0035 LUCI-2 #0390-0400

UT 11:09 - Sending Full AO w/N3.75 (R=9.43)

UT 11:12 - DX picked the wrong guidestar, also had to manually abort wait for collimation on DX (canceling preset was not enough).

UT 11:15 - sent again - the instrument config panels on each side of LUCI Observer GUI shows red, but preset is OK, guiding OK, AO REF ok - bug in software?

UT 11:25 - right side is OK, left side having issues - setting to psuedo-monocular DX , turning off readout/instrument/mirror for SX wil attempt to continue - didn't work, Geno clearing sync preset

UT 11:31 - made a L2 only script but got an error during commit - failed again

UT 11:32 - reconfigure for LUCI-2 only - cleared out stuck offset on left

UT 11:35 - got an error immediately on the instrument config - restarting scheduler on LUCI

UT 11:37 - error during commit - perhaps this is related to using twin.py?

UT 11:38 - taking data on DX only - DIMM 0.7" ~30 minutes left to sunrise - going to do 3 on-images with NCPA off (first 3 on) and 3 with NCPA on

UT 11:55 - LUCI-2 0408- attempting to turn on NCPA, but integration has started - can't turn on NCPA because too bright and it will melt... (paraphrasing Doug)

UT 11:57 - WE ARE DONE ON SKY - putting in blinds 0401-0408 LUCI-2 flexure compensation was off - this affected our attempt for Full AO.

================================================================================================

UT 12:00 - Closing enclosure..... will do MODS and LUCI cals (need K flats for LUCI ESM tests)

UT 12:17 - LUCI cals look off, lights off have lots of counts (persistence?) - LUCI-2 #0419-0421 test exposure 2.51x3 to check counts (forgot to do integrated)

UT 12:18 - lamp off 16k counts in K 2.51sec, taking 2nd test exposure 0422 - DO NOT USE CALS LUCI-1 0036-0045 LUCI-2 0408-0426

UT 12:21 - waiting a bit for possible persistence to die down....no its thermal, put in J-band filter got 10-20 counts in 2.51 sec

UT 12:30 - Retaking ESM K-band flats LUCI-1 0045-0049 Lights Off, 0050-0054 Lights On; LUCI-2 0427-0431 lights off, 0432-0436 lights on

UT 12:31 - MODS cals proceeding...pixel flats, arcs, 1" slit flats

UT 12:34 - Putting LUCIs in safe config....hearing the siren call of bagels....

==============================================
Topic revision: r9 - 17 May 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback